OpenStudy (andriod09):

Find 3 consecutive even whole numbers that add to 2112.

OpenStudy (andriod09):

@cshalvey

OpenStudy (anonymous):

consider the numbers as x , x+2 , x+4 sum of all three number = 2112

OpenStudy (andriod09):

THank you.

OpenStudy (anonymous):

welcome.

OpenStudy (anonymous):

x+ x+2 + x +4 = 2112 3x+6=2112 3x=2112-6 3x=2106 x=2106/3 x=702 so x+2=702+2=704 and x+4=702+4=706 check: 702+704+706=2112 therefore the 3 consecutive numbers that total 2112 are 702,704,706